OpenStudy (anonymous):

can any1 help me in physics please

OpenStudy (anonymous):

prove \[s=ut+\frac{1}{2}at^2\]

OpenStudy (anonymous):

i think i can wait.....

OpenStudy (anonymous):

\[average velocity=(v+u) \div2\] \[distance covered,s=average velo timestime\] =\[[(v+u)\div2]timest\] =\[[(u+at+u)\div2]timest\] =\[(2ut+at ^{2})\div2\] =\[ut+(1\div2)at ^{2}\] therefore s=\[ut+(1\div2)at ^{2}\] hope this helps sorry that i made u wait:)

OpenStudy (anonymous):

times means 'x'

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

thanks